OpenStudy (anonymous):

Convert the equation to polar form. x^2-y^2=7

myininaya (myininaya):

x=rcos(theta) y=rsin(theta) (rcostheta)^2-(rsintheta)^2=7

OpenStudy (anonymous):

what is r and theta, that's what confuses me

myininaya (myininaya):

do you not see what I scanned for you?

OpenStudy (anonymous):

i did

myininaya (myininaya):

i put a point on (x,y) and theta is the angle between the initial ray and the vector with endpoint (0,0) and (x,y) ( r is like the length of this vector except it can be negative) r is like the radius of a circle sortof except it can be negative r is the directed distance from the origin so if is negative you go the opposite a way anyways i drew a triangle, a right trinagle and i found cos(theta)=x/r and sin(theta)=y/r so that means rcos(theta)=x and rsin(theta)=y

myininaya (myininaya):

so (x,y)=(rcos(theta),rsin(theta))

OpenStudy (anonymous):

i understand that but i have no idea when it comes to plugging in the numbers